OpenStudy (elenathehomeschooler):

can someone help me solve for x e^x e^(x+1)=1

MsBrains (ms-brains):

Is this all of your question?

OpenStudy (elenathehomeschooler):

\[e^x e ^{x+1} =1\]

OpenStudy (elenathehomeschooler):

yes

MsBrains (ms-brains):

What do you think it is?

zepdrix (zepdrix):

Recall your exponent rule:\[\large\rm a^b\cdot a^c\quad=a^{b+c}\]

zepdrix (zepdrix):

Do you see how that will help us on the left side of the equation?

OpenStudy (elenathehomeschooler):

yes

OpenStudy (elenathehomeschooler):

i think it might be 0

zepdrix (zepdrix):

For x? Hmm no.

zepdrix (zepdrix):

Applying the rule to the left side gives us:\[\large\rm e^{x+(x+1)}=1\]\[\large\rm e^{2x+1}=1\]Then apply natural log to each side. Or rewrite 1 as e^0. Whichever method you prefer

zepdrix (zepdrix):

\[\large\rm e^{2x+1}=e^0\]

OpenStudy (elenathehomeschooler):

would we cancel out e

zepdrix (zepdrix):

Bases are equivalent, both e, so the exponents must be equivalent as well, ya? :) \[\large\rm 2x+1=0\] Sure, you can think of it as cancelling out, ya

OpenStudy (elenathehomeschooler):

would we get -0.5

zepdrix (zepdrix):

Yes :)

OpenStudy (elenathehomeschooler):

yay! so the answer is x=-0.5

zepdrix (zepdrix):

yay good job

OpenStudy (elenathehomeschooler):

Thank you so much!